40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Feb 20 19:09:48 2015, ericq, Update, ASC, QPD frontend code unified 
    Reply  Mon Feb 23 08:06:59 2015, rana, Update, ASC, QPD frontend code unified 
    Reply  Thu Mar 5 20:44:30 2015, Jenne, Update, SUS, damprestore script updated 
    Reply  Mon Mar 9 18:10:59 2015, Jenne, Update, ASC, Oplev filters re-copied 
    Reply  Sat Mar 14 00:12:18 2015, rana, Update, SUS, Oplevs huh? burp.png
       Reply  Mon Mar 23 13:27:33 2015, ericq, Update, SUS, ITMX oplev quadrant gains unbalanced olsum.png
          Reply  Fri Apr 3 14:57:38 2015, manasa, Update, SUS, BS oplev BS_oplev_Apr3.png
             Reply  Fri Apr 3 15:15:55 2015, Steve, Update, SUS, BS oplev 
                Reply  Fri Apr 3 19:35:14 2015, Jenne, Update, SUS, BS oplev centered 
Message ID: 11056     Entry time: Fri Feb 20 19:09:48 2015     Reply to this: 11057   11104   11125   11142
Author: ericq 
Type: Update 
Category: ASC 
Subject: QPD frontend code unified 

I have changed all of the oplevs and transmon QPDs to use the common ISC QPD library block, which differs mainly in its divide by zero protection. 

c1scx.mdl and c1scy.mdl were directly changed for the transmon QPDs. The oplevs were done by changing the sus_single_control.mdl library part, which is used for all of the SOSs. 

Then, because of the underscore introduced (i.e. OLPIT becomes OL_PIT because there is an OL block), I went on a sed safari to find and replace the new channel names into:

  • The filter ini files
  • various MEDM screens
  • The optic misaligning scripts (which currently live in medm/MISC/ifoalign, and need to get moved to scripts/)
  • A recent BURT snapshot, to restore all of the switches and settings easily. 
  • scripts/activateDQ.py, which is responsible for renaming OL_PIT_IN1_DQ to OPLEV_PERROR, etc.

I've fixed everything that occured to me, and the usual ways I'm used to interacting with the oplevs all seem to work at this time, but it's entirely possible I've overlooked something.

One important note is: because we are now using an effectively immutable QPD library block, the oplev urad conversion has to take place in the DoF matrix. The EPICS records C1:SUS-[OPTIC]_OL_[DOF]_CALIB still exist, but do not multiply the fast signals. Rather, the OL_MTRX elements are multiples of the CALIB value. I thought about making a new QPD_CALIBRATED part or something, but then we're right back to using custom code, which is what we're trying to avoid. 

All of the oplev DoFs are stable, I checked a few loop TFs like ETMY pitch and PRM yaw, and they looked normal. 

ELOG V3.1.3-